Evaluate \int \:\int _D\left(x+y\right)^{-\frac{1}{2}}dA\: over the region π‘₯ βˆ’ 2𝑦 ≀ 1 and π‘₯ β‰₯ 𝑦2 + 1


1
Expert's answer
2021-02-24T07:34:52-0500

Evaluate "\iint _D\left(x+y\right)^{-\frac{1}{2}}dA" : over the region π‘₯ βˆ’ 2𝑦 ≀ 1 and π‘₯ β‰₯ 𝑦2 + 1

Solution:

"\begin{cases}
 x-2y=1 \\
 x=y^2+1
\end{cases}" "\begin{cases}
 x=2y+1 \\
 x=y^2+1
\end{cases}" "\begin{cases}
 x=1 & y=0 \\
 x=5 & y=2
\end{cases}"

Region:

"\iint _D\left(x+y\right)^{-\frac{1}{2}}dA=\int_0^2dy\int_{y^2+1}^{2y+1}\frac{1}{\sqrt{x+y}}dx="

"2\int_0^2(\sqrt{3y+1}-\sqrt{y^2+y+1})dy="

"2\int_0^2(\sqrt{3y+1}-\sqrt{(y+\frac12)^2+\frac34})dy="

"\left(\frac49(3y+1)^{\frac32}-(y+\frac12)\sqrt{(y+\frac12)^2+\frac34}-\frac 34\ln{(y+\frac12+\sqrt{(y+\frac12)^2+\frac34})}\right)|_0^2\approx"

0.747874.

Answer: "\iint _D\left(x+y\right)^{-\frac{1}{2}}dA\approx0.747874" .